Dear tutors:
I have 3 problems due this evening involving using set builder notations
1) x+2 is greater than/equal to -3/11
2) -5/6x > 4/11
3) -10 (x+1) is less than/equal to 12x+7
Any help is appreciated.
kathrynneconway@yahoo.com

Answer by nerdybill(7384)

1) x+2 is greater than/equal to -3/11
.
x+2 >= -3/11
x >= -3/11 - 2
x >= -3/11 - 22/11
x >= -25/11
{x : x >= -25/11}
.
2) -5/6x > 4/11
-5/6x > 4/11
-5x > 24/11
-x > 24/55
x < -24/55
{x : x < -24/55}
.
3) -10 (x+1) is less than/equal to 12x+7
-10(x+1) <= 12x+7
-10x-10 <= 12x+7
-10 <= 22x+7
-17 <= 22x
-17/22 <= x
{x : x >= -17/22}
.